# Wilsonian Renormalization of Neural Network Gaussian Processes

### v1

@@ 4 Renormalization beyond the Gaussian
$Z^{M}\right>_{\eta}=e^{-\eta}\int\!\mathcal{D}\bm{f}_{<}\,e^{-S_{0}[\bm{f}_{<}]}\int\!\mathcal{D}\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]+\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]}\leavevmode\nobreak\ .$ (49)
@@
$\int\!\mathcal{D}$ $\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]+\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,\exp\left[-\frac{1}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}+\frac{1+2B}{2\sigma^{4}}\left(\Phi_{<}^{\top}\bm{f}_{q}\right)^{2}-\frac{A}{\sigma^{2}}\,\Phi_{<}^{\top}\bm{f}_{q}\right]}$ (50)
$\approx$ $\exp\left\{\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{2}}+\lambda_{q}(1+2B)\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{4}}}\right\}$
@@
4.1 A solvable toy example: single mode integration
@@
As a concrete application, here we consider perhaps the simplest setting with tractable non-Gaussian effects. However, unlike in the above examples where such effects were treated perturbatively, here we consider a more natural setting relevant to fully connected networks, wherein the lesser feature modes largely determine the greater feature modes. Concretely, we consider $x\in\mathbb{R}$ with $\mathrm{d}\mu_{x}\sim{\cal N}[0,1]$, and a rank-2 kernel of the form
@@
in particular, $He_{1}(x)=x$ and $He_{2}(x)=x^{2}-1$. Given the Gaussian measure, $He_{1}(x)$, $He_{2}(x)$ can be shown to span the non-zero eigenspaces of the kernel with eigenvalues ${\lambda_{1}:=\lambda_{<}}$, ${\lambda_{2}:=\lambda_{>}}$ respectively. To connect with our formalism above, note that here we have only two modes, $k\in\{1,2\}$, which are furnished by the first two Hermite polynomials, i.e., $\phi_{1}(x)=He_{1}(x)$ and $\phi_{2}(x)=He_{2}(x)$, cf. (7),
Our task is to learn a target function
@@
From a spectral bias perspective canatar , the average predictor for this model is zero, since the target is orthogonal to all non-zero eigenspaces under the input measure. [^11] By this we mean that the inner product on the associated Hilbert space vanishes, $\int\!\mathrm{d}\mu_{x}\,K(x,y)He_{5}(x)=0$. This is not necessarily the case however, since the features ($\varphi$, defined in (20)) will be non-Gaussian unless $\phi(x)$ is an affine function of $x\sim\mathcal{N}[0,1]$. Statistically therefore, the more common events in which $x$ lies closer to zero may bias the predictor on small datasets to act as if it were seeing the function $-10x^{3}+15x$, which does have non-vanishing overlap with $He_{1}(x)$. We show below that our RG approach can accurately predict such effects, and that they can be viewed as stemming from an MSE loss weighted by $\left(1-\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)$. Indeed, while $y=He_{5}(x)$ and $He_{1}(x)$ are not coupled by standard MSE loss, the extra weighting term yields a non-zero overlap via integrals of the type $\int\!\mathrm{d}\mu_{x}He_{1}(x)\left(\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)He_{5}(x)$.
We shall first derive the above weighted form of the MSE loss, and then discuss its effect on predictions. Since the kernel mode associated with $He_{5}(x)$ is frozen/non-fluctuating, we do not treat it as a GP mode, and instead integrate out only the mode associated with $\lambda_{2}$. [^12] That is, we view $\bm{y}_{<}(x)$ as $\bm{y}_{1}He_{1}(x)+\bm{y}_{5}He_{5}(x)=He_{5}(x)$, with $y_{1}=0$ and $y_{5}=1$, and similarly $f_{<}(x)=f_{1}x$ with $f_{5}=0$. Alternatively, one may treat $y$ as part of the greater target modes, where its effect would then enter as a subleading shift to the lesser target. We denote this by $\varphi_{>}=\varphi_{2}$, and the associated replica-space GP mode by $\bm{f}_{2}$, with $\bm{y}_{2}=0$ (i.e., $q=2$ in our notation above). Furthermore, we take ${\bm{f}}_{1}$ to be the replica space vector associated with the linear (lesser) mode and (i.e., $He_{1}(x)$) and ${\bm{y}}_{1}$ to be the corresponding target. Converting to feature variables $\varphi_{<}=\varphi_{1}$ and $\varphi_{>}=\varphi_{2}$ via (20), we have
@@
$P[\varphi_{1},\varphi_{2}]$ $=P[\varphi_{1}]P[\varphi_{2}|\varphi_{1}]=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\int\!\mathrm{d}\mu_{x}\delta\left(\varphi_{2}-He_{2}(x)\right)$
$=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\delta\left(\varphi_{2}-He_{2}(\varphi_{1})\right)=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\delta\left[\varphi_{2}-\left(\varphi_{1}^{2}-1\right)\right]\leavevmode\nobreak\ .$
@@
$=-\eta\int\!\mathrm{d}\varphi_{1}\,P[\varphi_{1}]\exp\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{1}^{\top}\Phi_{1}+He_{2}(\varphi_{1})^{2}|\bm{f}_{2}|^{2}+2He_{2}(\varphi_{1})He_{1}(\varphi_{1})\bm{f}_{2}^{\top}(\bm{f}_{1}-\bm{y}_{1})\right)\right\}$
@@
$Z^{M}\right>_{\eta}$ $=\int\!\mathcal{D}f_{1}\exp\left\{-S_{0}[\bm{f}_{1}]+\eta\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(\varphi_{1})^{2}\right)}\right\}$ (60)
$=\int\!\mathcal{D}f_{1}\exp\left\{-S_{0}[\bm{f}_{1}]+\eta\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(x)^{2}\right)}\right\}\leavevmode\nobreak\ ,$
where on the second line we used (58) to replaced $\varphi_{1}$, $\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]$ with $x$, $\int\!\mathrm{d}\mu_{x}$, respectively. We thus observe the advertised reweighting of the MSE term, namely
@@
Next, we perform an EK-type approximation on (60). Namely, we assume $\sigma^{2}$ is large enough compared to $\left(\bm{f}_{1}(x)-\bm{y}_{1}(x)\right)$ to perform a Taylor expansion of the exponent appearing in the action. We note in passing that the next-to-leading order in this expansion Cohen_2021 . happens to vanish for $He_{n\geq 5}(x)$ (but not, e.g., for $He_{3}(x)$). Within the leading order expansion, the action in (60) becomes
$\frac{1}{2\lambda_{1}}\bm{f}_{1}^{2}-\eta\int\mathrm{d}\mu_{x}\left(1-\frac{1}{2\sigma^{2}}(\bm{f}_{1}x-He_{5}(x))^{2}\left[1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(x)^{2}\right]\right)\leavevmode\nobreak\ .$ (62)
@@
$\bar{f}_{1}\approx-\frac{\lambda_{1}}{\lambda_{1}+\sigma^{2}/\eta}\frac{120\lambda_{2}}{\sigma^{2}}\leavevmode\nobreak\ .$ (63)
@@
To test this empirically, we sampled various draws of datasets with $n=100$ from $x\sim{\cal N}[0,1]$, performed GP regression on these with the above kernel using a fixed random test set with $n_{\mathrm{test}}=1000$, and averaged the results over the various draws. Doing so for $\lambda_{2}=0$ shows an almost zero outcome, as expected. However, taking finite $\lambda_{2}$ reveals a trend that spectral bias Cohen_2021 ; canatar misses. Notably, the fact that our toy example involves only one mode which is integrated out, together with the need to take its variance to be small ($\lambda_{2}/\sigma^{2}\ll 1$) as well as make it unlearnable ($n\lambda_{2}/\sigma^{2}\ll 1$) results in a rather small effect. Still, it serves to demonstrate the conceptual phenomena whose importance depends on the setting or the desired accuracy.
@@
Figure 1: Non-Gaussian features and spatial re-weighting effects. Theory versus experiment for the model of sec. 4.1 with $n=100$ datapoints $\sigma^{2}=400$ and $\lambda_{>}:=\lambda_{2}=0.1$ (unless stated otherwise). Learning a 5th Hermite polynomial, using a kernel capable of expressing only 1st and 2nd Hermite polynomials should give a zero average predictor (green line) based on the standard theory canatar ; Cohen_2021 . However, due to spatial re-weighing, a coupling between 1st and 5th Hermite polynomial arises leading to a non-zero result.
5 Discussion

### v2

@@ IV Renormalization beyond the Gaussian
$Z^{M}\right>_{\eta}$ $=e^{-\eta}\int\!\mathcal{D}\bm{f}_{<}\,e^{-S_{0}[\bm{f}_{<}]}\int\!\mathcal{D}\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]-S_{\mathrm{int}}}\leavevmode\nobreak\ .\quad$ (54)
@@
$\int\!\mathcal{D}\bm{f}_{q}\,e^{-S_{0}[\bm{f}_{q}]-S_{\mathrm{int}}}$
$\approx\exp\left\{\eta\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\,e^{-\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{2}}+\lambda_{q}(1+2B)\frac{\Phi_{<}^{\top}\Phi_{<}}{2\sigma^{4}}}\right\}\leavevmode\nobreak\ ,\quad$ (55)
@@
IV.1 Functional RG flow
Extending the above single-mode integration into a potential flow equation requires handling two complications. The first is the need to express the distribution of features $P[\varphi_{<}]$ which was so far left abstract. To this end, we should marginalize over $\phi_{q}$ in $P[\varphi]$. Conveniently, within the Edgeworth expansion, this marginalization amounts to omitting the $\varphi_{q}$ mode. This can be seen from (37) and the fact that $He_{4}$ with any number of fixed lesser arguments integrates to zero under the greater Gaussian distribution. For instance, given one lesser argument and three greater arguments, we have
$\int$ $\mathrm{d}\varphi_{q}\,\mathcal{N}[0,1;\varphi_{q}]He_{4}[\varphi_{k},\varphi_{q},\varphi_{q},\varphi_{q}]$ (60)
$=\int\!\mathrm{d}\varphi_{q}\,\mathcal{N}[0,I;\varphi_{k}]^{-1}\partial_{\varphi_{k}}\mathcal{N}[0,I;\varphi_{k}]\,\partial_{\varphi_{q}}^{3}\mathcal{N}[0,I;\varphi_{q}]$
$=\varphi_{k}\!\int\!\mathrm{d}\varphi_{q}\left(\varphi_{q}^{3}-3\varphi_{q}\right)\mathcal{N}[0,I;\varphi_{q}]=0\leavevmode\nobreak\ .$
The second issue is the appearance of a new structure – the reweighting of the $L^{2}$ norm – in the one-step renormalized action, which necessitates its inclusion in further RG steps. To this end, we repeat the previous derivation in its presence but take $y_{>}(x)=0$ for simplicity. First, we write $W(x)=1+\delta W(x)$, and take the following ansatz for the infinitesimal change
$\delta W(x)=\sum_{k_{1},k_{2}}W_{k_{1}k_{2}}He_{2}[\phi_{k_{1}}(x),\phi_{k_{2}}(x)]\leavevmode\nobreak\ ,$ (61)
where based on (57) we have that at the second RG step we are currently carrying $W_{k_{1}k_{2}}=-\frac{\delta c}{\sigma_{0}^{2}}U_{k_{1}k_{2}qq}$. We shall see below (68) that our RG flow respects this ansatz. Notably, for the current RG step, we denote the greater mode by $q^{\prime}$. As $\delta W(x)$ contains pairs of features ($\varphi$’s) we split it into its lesser and greater components
$\delta W_{<}(x)$ $=\sum_{k_{1},k_{2}\neq q^{\prime}}W_{k_{1}k_{2}}He_{2}[\varphi_{k_{1}},\varphi_{k_{2}}]\leavevmode\nobreak\ ,$ (62)
$\delta W_{>}(x)$ $=2\sum_{k_{1}\neq q^{\prime}}W_{k_{1}q^{\prime}}He_{2}[\varphi_{k_{1}},\varphi_{q^{\prime}}]$ (63)
$\qquad\qquad\qquad+W_{q^{\prime},q^{\prime}}He_{2}[\varphi_{q^{\prime}},\varphi_{q^{\prime}}]\leavevmode\nobreak\ .$
Following this, we may decompose the interaction term into lesser/greater components as
$S$ int
$=$ $\int\!\mathcal{D}\varphi\,P[\varphi_{<},\varphi_{>}]\,$
×$\times\exp\!\left\{-\frac{1+\delta W_{<}+\delta W_{>}}{2\sigma^{2}}\left(\Phi_{<}^{\top}\Phi_{<}+\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right\}\leavevmode\nobreak$
$=$ $\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]$ (64)
×$\times\exp\!\left\{-\frac{1+\delta W_{<}+\delta W_{>}}{2\sigma^{2}}\left(\Phi_{>}^{\top}\Phi_{>}+2\Phi_{<}^{\top}\Phi_{>}\right)\right.$
$\qquad\qquad\qquad\qquad\left.-\frac{1}{2\sigma^{2}}\delta W_{>}\Phi_{<}^{\top}\Phi_{<}\right\}\leavevmode\nobreak\ ,$
cf. (53).
Consistent with our earlier derivation, we aim to keep the leading order contribution in $U$, the non-Gaussian correction. We may thus expand the above to linear order in $\delta W_{\lessgtr}$. As before (cf. (46)), we also drop the $\Phi_{>}^{\top}\Phi_{>}$ term, thus leading to
$S_{\mathrm{int}}=$ $\int\!\mathcal{D}\varphi_{<}\,P[\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right\}$ (65)
×$\times\int\!\mathcal{D}\varphi_{>}P[\varphi_{>}|\varphi_{<}]\exp\!\left\{-\frac{1+\delta W_{<}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right\}\leavevmode\nobreak$
×$\times\left(1-\frac{\delta W_{>}}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}-\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right)\leavevmode\nobreak\ .$
We next observe that if we take $\delta W_{>}=0$, the previous derivation could be repeated as-is under the replacement $\frac{1}{\sigma^{2}}\rightarrow\frac{1+\delta W_{<}}{\sigma^{2}}$. This is because $\delta W_{<}$ is a constant with respect to both the greater feature mode integration and the greater GP modes, and thus acts as a parameter under all algebraic manipulations carried out in the previous section.
Our emphasis thus turns to the two $\delta W_{>}$ dependent terms. As these are linear order in $U$ (the non-Gaussian correction), under our leading order in $U$ treatment, we may evaluate their contribution while taking $\delta W_{<}=0$ (since $\delta W$ is $O(U)$) and $P[\varphi_{>}|\varphi_{<}]={\cal N}[0,I;\varphi_{>}]$ (since corrections to Gaussianity are $O(U)$). Namely, we may simplify the $\delta W_{>}$ dependence of the $\int\!\mathcal{D}\varphi_{>}$ integral into
$-\int\!\mathcal{D}\varphi_{>}{\cal N}[0,1;\varphi_{>}]\left(1-\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{1}{2}\left[\frac{1}{\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}\right]^{2}\right)$
×$\qquad\times\left(\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{>}+\frac{\delta W_{>}}{2\sigma^{2}}\Phi_{<}^{\top}\Phi_{<}\right),$ (66)
where in the first bracket we kept up to second order in the greater GP mode since, as seen before, higher terms would contribute as $\delta c$ to a power larger than $1$ following the greater GP mode integration. In the second bracket, we kept leading order in $U$.
The above integral leads to five terms, whose contributions to the renormalized GP action are analyzed in appendix C. The crucial observation that each such contribution appears as $O(\delta W_{>})O(\delta c/\sigma^{2})$ and consequently, as $\delta W_{>}$ is itself $O(\delta c/\sigma^{2})$— of order $(\delta c/\sigma^{2})^{2}$. Such contributions are negligible within our momentum-shell (i.e., leading order in $\delta c/\sigma^{2}$) treatment. Indeed $O(\delta c/\sigma^{2})^{2}$ terms, even after being summed $c/\delta c$-times during the RG process (where $c$ is the total variance of the integrated out modes), would be $O(c/\delta c)O(\delta c^{2})=O(\delta c)$ and vanish in the $\delta c\rightarrow 0$ limit. In contrast, the $O(\delta c)$ terms (“Wilsonian” terms) may contribute $O(c/\delta c)O(\delta c)=O(c)$ and hence remain finite as $\delta c\rightarrow 0$.
The above considerations imply that we may ignore all greater components of the weighting term in the RG process, and assume $W(x)$ contains only lesser feature modes. With this major simplification, we can proceed as before, with $\Phi_{<}^{T}\Phi_{>}/\sigma_{c}^{2}$ replaced by $W_{<}\Phi_{<}^{T}\Phi_{>}/\sigma_{c}^{2}$. As mentioned above, since $W_{<}$ is constant with respect to all RG manipulations (i.e., it is independent of the greater feature modes and all GP modes), this amounts to just carrying it through as a factor accompanying $\sigma^{-2}$. We thus find
$\frac{W_{2\delta c}(x)}{\sigma_{2\delta c}^{2}}=\frac{W_{\delta c}(x)}{\sigma_{\delta c}^{2}}-\frac{\delta cW_{\delta c}(x)^{2}(1+2B_{\delta c}(x))}{\sigma^{4}_{\delta c}}$
$=\frac{1-\frac{2\,\delta c}{\sigma^{2}_{0}}B_{0}(x)}{\sigma_{\delta c}^{2}}-\frac{\delta c\left(1-\frac{2\,\delta c}{\sigma^{2}_{0}}B_{0}(x)\right)^{2}(1+2B_{\delta c}(x))}{\sigma^{4}_{\delta c}}$
$=\frac{1-\frac{2\,\delta c}{\sigma^{2}_{0}}B_{0}(x)}{\sigma_{\delta c}^{2}}-\frac{\delta c(1+2B_{\delta c}(x))}{\sigma^{4}_{\delta c}}+O(\delta c^{2})\leavevmode\nobreak\ .$ (67)
As before, integrating both sides with respect to $\mathrm{d}\mu_{x}$ recovers the previous RG flow of $\sigma^{2}_{2\delta c}$, namely $\sigma^{2}_{2\delta c}=\sigma^{2}_{\delta c}+\delta c$ (we assume without loss of generality that both integrated modes have the same $\delta c$). Following a similar algebra to the our first RG step described in the previous subsection, we find
$W_{2\delta c}(x)$ $=W_{\delta c}(x)-\frac{2\,\delta c}{\sigma_{\delta c}^{2}}[B_{\delta c}(x)]+O(\delta c^{2})$ (68)
$=W_{0}(x)-\frac{2\,\delta c}{\sigma_{\delta c}^{2}}[B_{\delta c}(x)+B_{0}(x)]\leavevmode\nobreak\ ,$
which again integrates to 1 under $\mathrm{d}\mu_{x}$ (i.e., normalization is preserved).
This time, no new structure appeared in our RG step, hence we may iterate it $M$ times to obtain the following RG flow
$\sigma^{2}_{c}$ $=\sigma^{2}_{0}+c$ (69)
$W_{c}(x)$ $=1-\sum_{m=0}^{M-1}\delta cB_{m\delta c}(x)/\sigma_{c}$ (70)
[…]
where $\delta cM=c$ and the integral is obtain in the limit $\delta c\rightarrow 0$.
IV.2 A solvable toy example: single mode integration
@@
As a concrete application, here we consider perhaps the simplest setting with tractable non-Gaussian effects. However, unlike in the above examples where such effects were treated perturbatively, here we consider a natural setting relevant to fully connected networks, wherein the lesser feature modes largely determine the greater feature modes. Concretely, we consider $x\in\mathbb{R}$ with $\mathrm{d}\mu_{x}\sim{\cal N}[0,1]$, and a rank-2 kernel of the form
@@
In particular, $He_{1}(x)=x$ and $He_{2}(x)=x^{2}-1$. Given the Gaussian measure, $He_{1}(x)$, $He_{2}(x)$ can be shown to span the non-zero eigenspaces of the kernel with eigenvalues ${\lambda_{1}:=\lambda_{<}}$, ${\lambda_{2}:=\lambda_{>}}$ respectively. To connect with our formalism above, note that here we have only two modes, $k\in\{1,2\}$, which are furnished by the first two Hermite polynomials, i.e., $\phi_{1}(x)=He_{1}(x)$ and $\phi_{2}(x)=He_{2}(x)$, cf. (7).
Our task is to learn a target function,
@@
From a spectral bias perspective [5], the average predictor for this model is zero, since the target is orthogonal to all non-zero eigenspaces under the input measure. [^11] By this we mean that the inner product on the associated Hilbert space vanishes, $\int\!\mathrm{d}\mu_{x}\,K(x,y)He_{5}(x)=0$. This is not necessarily the case however, since the features ($\varphi$, defined in (20)) will be non-Gaussian unless $\phi(x)$ is an affine function of $x\sim\mathcal{N}[0,1]$. Statistically therefore, the more common events in which $x$ lies closer to zero may bias the predictor on small datasets to act as if it were seeing the function $-10x^{3}+15x$, which does have non-vanishing overlap with $He_{1}(x)$. We show below that our RG approach can accurately predict such effects, and that they can be viewed as stemming from an MSE loss weighted by $\left(1-\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)$. Indeed, while $y=He_{5}(x)$ and $He_{1}(x)$ are not coupled by standard MSE loss, the extra weighting term yields a non-zero overlap via integrals of the type $\int\!\mathrm{d}\mu_{x}He_{1}(x)\left(\lambda_{2}\sigma^{-2}He_{2}(x)^{2}\right)He_{5}(x)$.
We shall first derive the above weighted form of the MSE loss, and then discuss its effect on predictions. Since the kernel mode associated with $He_{5}(x)$ is frozen/non-fluctuating, we do not treat it as a GP mode, and instead integrate out only the mode associated with $\lambda_{2}$. [^12] That is, we view $\bm{y}_{<}(x)$ as $\bm{y}_{1}He_{1}(x)+\bm{y}_{5}He_{5}(x)=He_{5}(x)$, with $y_{1}=0$ and $y_{5}=1$, and similarly $f_{<}(x)=f_{1}x$ with $f_{5}=0$. Alternatively, one may treat $y$ as part of the greater target modes, where its effect would then enter as a subleading shift to the lesser target. We denote this by $\varphi_{>}\eqqcolon\varphi_{2}$, and the associated replica-space GP mode by $\bm{f}_{2}$, with $\bm{y}_{2}=0$ (i.e., $q=2$ in our notation above). Furthermore, we take ${\bm{f}}_{1}$ to be the replica space vector associated with the linear (lesser) mode and (i.e., $He_{1}(x)$) and ${\bm{y}}_{1}$ to be the corresponding target. Converting to feature variables $\varphi_{<}\eqqcolon\varphi_{1}$ and $\varphi_{>}\eqqcolon\varphi_{2}$ via (22), we have
@@
$P[\varphi_{1},\varphi_{2}]$ $=P[\varphi_{1}]P[\varphi_{2}|\varphi_{1}]$ (75)
$=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\int\!\mathrm{d}\mu_{x}\delta\left(\varphi_{2}-He_{2}(x)\right)$
$=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\delta\left(\varphi_{2}-He_{2}(\varphi_{1})\right)$
$=\mathcal{N}[0,\mathbb{1};\varphi_{1}]\delta\left[\varphi_{2}-\left(\varphi_{1}^{2}-1\right)\right]\leavevmode\nobreak\ .$
@@
$=-\eta\int\!\mathrm{d}\varphi_{1}\,P[\varphi_{1}]\exp\left\{-\frac{1}{2\sigma^{2}}\left(\Phi_{1}^{\top}\Phi_{1}+He_{2}(\varphi_{1})^{2}|\bm{f}_{2}|^{2}+2He_{2}(\varphi_{1})He_{1}(\varphi_{1})\bm{f}_{2}^{\top}(\bm{f}_{1}-\bm{y}_{1})\right)\right\}.$
@@
$Z^{M}\right>_{\eta}$ $=\int\!\mathcal{D}f_{1}\exp\bigg{\{}-S_{0}[\bm{f}_{1}]$ (77)
$\qquad+\eta\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(\varphi_{1})^{2}\right)}\bigg{\}}$
$=\int\!\mathcal{D}f_{1}\exp\bigg{\{}-S_{0}[\bm{f}_{1}]$
$\qquad+\eta\int\!\mathrm{d}\mu_{x}\,e^{-\frac{1}{2\sigma^{2}}\Phi_{1}^{\top}\Phi_{1}\left(1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(x)^{2}\right)}\bigg{\}}\leavevmode\nobreak\ ,$
where on the second line we used (75) to replace $\varphi_{1}$, $\int\!\mathrm{d}\varphi_{1}P[\varphi_{1}]$ with $x$, $\int\!\mathrm{d}\mu_{x}$, respectively. We thus observe the advertised reweighting of the MSE term, namely
@@
Next, we perform an EK-type approximation on (77). Namely, we assume $\sigma^{2}$ is large enough compared to $\left(\bm{f}_{1}(x)-\bm{y}_{1}(x)\right)$ to perform a Taylor expansion of the exponent appearing in the action. We note in passing that the next-to-leading order in this expansion [32] happens to vanish for $He_{n\geq 5}(x)$ (but not, e.g., for $He_{3}(x)$). Within the leading order expansion, the action in (77) becomes
$S_{0}[\bm{f}_{1}]+S_{\mathrm{int}}=\frac{1}{2\lambda_{1}}\bm{f}_{1}^{2}-\eta\int\mathrm{d}\mu_{x}\left(1-\frac{1}{2\sigma^{2}}(\bm{f}_{1}x-He_{5}(x))^{2}\left[1-\frac{\lambda_{2}}{\sigma^{2}}He_{2}(x)^{2}\right]\right)\leavevmode\nobreak\ .$ (79)
@@
$\bar{f}_{1}\approx-\frac{\lambda_{1}}{\lambda_{1}+\sigma^{2}/\eta}\frac{120\lambda_{2}}{\sigma^{2}}\leavevmode\nobreak\ ,$ (80)
@@
To test this empirically, we sampled various draws of datasets with $n=100$ from $x\sim{\cal N}[0,1]$, performed GP regression on these with the above kernel using a fixed random test set with $n_{\mathrm{test}}=1000$, and averaged the results over the various draws (see fig. 1). Doing so for $\lambda_{2}=0$ shows an almost zero outcome, as expected. However, taking finite $\lambda_{2}$ reveals a trend that spectral bias [32, 5] misses. Notably, the fact that our toy example involves only one mode which is integrated out, together with the need to take its variance to be small ($\lambda_{2}/\sigma^{2}\ll 1$) as well as make it unlearnable ($\eta\lambda_{2}/\sigma^{2}\ll 1$) results in a rather small effect. Still, it serves to demonstrate the conceptual phenomena whose importance depends on the setting or the desired accuracy.
@@
Figure 1: Non-Gaussian features and spatial re-weighting effects. Theory versus experiment for the model of sec. IV.2 with $n=100$ datapoints $\sigma^{2}=400$ and $\lambda_{>}:=\lambda_{2}=0.1$ (unless stated otherwise). Learning a 5th Hermite polynomial, using a kernel capable of expressing only 1st and 2nd Hermite polynomials should give a zero average predictor (green line) based on the standard theory [5, 32]. However, due to spatial re-weighting, a coupling between 1st and 5th Hermite polynomial arises leading to a non-zero result.
V Discussion